OpenStudy (anonymous):

Find the length of 2v if v = <2, -4, 5>

OpenStudy (anonymous):

\[1. 3\sqrt{5}\] \[2. \sqrt{13}\] \[3. 6\sqrt{5}\] \[4. \sqrt{13\div2}\]